OpenStudy (anonymous):

Evaluate. Please show all of your work. 4 + x3 – x, for x = -3

OpenStudy (anonymous):

Substitute 3 in for x wherever it appears. \[4 + (-3)^3 - (-3)\]

OpenStudy (anonymous):

-3 sorry.

OpenStudy (anonymous):

x=-3 so replace all the x values with -3. and the solve like Above....

OpenStudy (anonymous):

4+(-3)*3-(-3) = 4+(-9)+3 = 7+(-9) = -2

OpenStudy (anonymous):

Unless you mean x^3 instead of x*3, then @qpHalcy0n is right.

OpenStudy (anonymous):

thank you thats the same answer i got i was just checking it